# Diffusion-VLA: Scaling Robot Foundation Models via Unified Diffusion and Autoregression

### v1

@@ 4 Experiments
Figure 4: Examples of visual variations, including randomly placed distractors, distracting lighting, and a colorful background on the right. DiVLA is robust to visual changes in different scenarios.
@@ 4.2 Real-World Multi-Task Learning
We begin with a standard setting in which the model is trained on multiple tasks and completes each task based on different user queries. We designed five tasks: object selection, flip the vertically placed pot, placing a cube into a designated box, placing a cup onto a plate, and placing a cube inside a box. Detailed descriptions of these tasks are provided in the Appendix. The experimental results can be found in Table 1.
We compare our method to the Diffusion Policy [10], TinyVLA [57], Octo [34], and OpenVLA [26]. Notice that both Octo and OpenVLA is pre-trained on OXE [35], which is 25 times larger than our pre-trained datasets.
@@ 4.3 End-to-End Sorting on Real Robot
We evaluate the capability of DiVLA in an industrial setting, where a robot is tasked with sorting items into designated sectors within a large box based on object category. Specifically, we categorize items into four classes: (1) toy cars, (2) knit gloves, (3) stuffed toys, and (4) hex keys. The language instruction provided is ”Sort all items into corresponding areas”. A total of 500 trajectories are collected as training data. The task is considered successful only if the robot successfully grasps the object and places it in the correct sector. The experimental setup is illustrated in Figure 1.
@@
Diagnosing policy model by checking reasoning. Because our model generates output with natural language reasoning, we can gain insight into what the model is “thinking” by observing its reasoning phrases. For example, as shown in Figure 5, the model identifies a toy car and decides to pick it up. If we intervene by placing a hex key in the gripper instead, the reasoning phrase shifts from “grabbing the toy car” to “grabbing the hex,” allowing the model to adapt and sort the item accurately. This dynamic reasoning makes the model’s decision-making process more transparent and interpretable. The reasoning injection module also benefits from the reasoning self-correction such that the robot action is more robust.
Identifing and sorting unseen objects. We observe that our model can identify unseen objects by name. When we place four previously unseen objects on the plate, a stuffed toy cat, a pair of green gloves, a dark toy car, and a screwdriver. The model successfully identifies the first two objects as “brown toy cat” and “green gloves”, and then correctly sorts them. For the remaining two objects on the right, which it cannot recognize, the model labels them as “the black object” and “the object”. Although it fails to sort these unrecognized items, our model still demonstrates generalization in identifying objects and effectively translates these identifications into appropriate actions.
Figure 5: DiVLA adjusts its output reasoning based on changes to the target object.
@@
Figure 7: Examples of various unseen objects in zero-shot bin-picking tasks. The unseen objects vary significantly in size, demonstrating that DiVLA  can successfully pick objects of different sizes without prior training on them.
Following Novel Instruction on Four Tasks
Model $\setminus$ Tasks Watermelon Watermelon $\rightarrow$ Blue Paper Trash $\rightarrow$ Lemonade
OpenVLA [26] 2/3 0/3
DiVLA-2B 2/3 2/3
[…]
OpenVLA [26] 0/3 0/3
DiVLA-2B 1/3 1/3
@@ 4.5 Adapt to Real-World Bimanual Robot
Figure 8: (a) Environmental setup for the bimanual robot, featuring three camera views; (b) Table bussing setup, with a trash bin positioned on the right side and a panel on the left. The task requires placing all tableware on the panel and all trash in the trash bin; (c-f) All tableware and trash items used in the bussing table task evaluation.
In this section, we examine DiVLA’s adaptability to a dual-arm robot. Inspired by $\pi_{0}$ [7], we designed a table bussing task that involves clearing table with various objects. This task was adapted for a bimanual robot setup: all tableware should be placed on a panel to the left, while trash items are dropped into a bin on the right. Similar to our factory sorting task, we evaluated the model’s performance using both seen objects and a combination of seen and unseen objects. The environment setup, along with all objects used for training and evaluation, is shown in Figure 8. Our evaluation consisted of twelve trials, with three to five objects randomly placed on the table. The success rate is computed by how many objects are correctly placed.
Our experimental results show that our model successfully completes tasks in most cases when the object has appeared in the training data, achieving an average success rate of 72.9% on seen objects. In contrast, the Diffusion Policy and OpenVLA achieve 45.8% and 0% success rates. For tasks involving both seen and unseen objects, DiVLA achieves up to a 70.8% success rate, a slight decrease in success rate compared to the seen one, showing remarkable generalization to objects with different colors and shapes. Finally, DiVLA demonstrates the ability to recognize unseen objects, particularly by responding sensitively to object color. For example, it categorizes a Sprite can as a ”green can” and correctly places it in the trash bin. This observation further supports the idea that reasoning contributes to generalization.
@@
4.6 Following Novel Instruction
The previous section primarily evaluated the model’s performance on in-distribution task commands. In this section, we assess the model’s ability to follow novel instructions, specifically focusing on its generalization to previously unseen commands. We introduce new instructions that prompt the model to
We tested the model on four objects: 1) a watermelon, 2) a lemonade, 3) a blue paper trash, and 4) a red pepper. This is an extremely challenging task, as these novel instructions are absent from both the Droid dataset and our collected data. We evaluated four new instructions, with results summarized in Table 2.
Our findings indicate that both OpenVLA and DiVLA-2B can recognize these unseen objects and perform basic pick-and-place tasks. However, when it comes to complex sequential tasks, OpenVLA fails to interpret the instructions accurately; instead, it randomly selects items. In contrast, our method correctly follows the instructions, picking up objects in the specified sequence. We hypothesize that by learning to decompose long tasks into subtasks, our method acquires a generalized ability to understand complex, multi-step instructions. While OpenVLA can execute simpler commands like ’Pick up the watermelon,’ it struggles with more advanced instructions requiring item selection in a specific order. Additionally, we observe a decrease in grasping precision when the model processes novel instructions, indicating that the novelty of instructions introduces further complexity to task execution.
@@ 5 Conclusion
In this work, we present DiVLA, a state-of-the-art vision-language-action model that delivers strong performance in both simulations and real-world scenarios, including single-arm and dual-arm robots. The core of our method lies in combining the next-token prediction objectives and diffusion models: the former for task reasoning and the latter for action prediction. We introduce a reasoning reuse module to enhance action generation and implement view-adaptive tokenization to reduce computational costs. Through extensive evaluations in both simulations and across multiple real-world embodiments, we demonstrate that DiVLA outperforms several SOTA robot models. Additionally, we show that DiVLA has robust generalization capabilities, adapting effectively to new instructions, tasks, and environments. Our research offers a novel perspective on designing VLA models, encouraging a rethinking of how reasoning can be reused to facilitate end-to-end policy learning.
@@
# Task # of Demonstrations Avg. Trajectory Length
1 Object selection 160 91
2 Flip the vertically placed pot 120 137
3 Place cube inside box with closed lid 50 146
4 Place cup onto the plate 100 107
5 Place cube into yellow/blue Box 100 90
@@ 6.1 Evaluation Tasks
As described in Section 4, we evaluate our DiVLA and baselines on multi-task learning, factory sorting, zero-shot bin-picking, and table bussing. In this section, we provide details on these tasks. All of these tasks are depicted in Figure 9 and 10.
For multi-task learning, we evaluate all methods on 5 tasks for multi-task learning and 5 visual generalization tasks. We evaluate each method with a total of 77 trials for multi-task learning and 45 trials for visual generalization. We provide the number of demonstrations and the average length of these tasks in 4. A brief description of these tasks is as follows:
@@
Moreover, detailed descriptions for factory sorting, zero-shot bin-picking, and table bussing are provided in Sections 4.3, 4.4, and 4.5, respectively.
@@ 6.2 Detailed Results
Task $\setminus$ Method Diffusion Policy OpenVLA DiVLA-2B
Sorting 0 0 60
@@ 7 Implementation Details
For the baselines, we follow a uniform training strategy to ensure consistency. For OpenVLA, we follow the original implementation which only uses a single camera. We also test OpenVLA with three camera views, where each view goes through the same visual encoder and concatenates its outputs for further process. We observe some performance improvement over multiple camera views. For instance, for sorting task, the average success rate will increase from 32.3% to 45.3%. Nevertheless, since most of the existing works only report the OpenVLA baseline with a single camera view, we only report experimental results for this setting. More details can refer to TinyVLA [57]. OpenVLA’s pre-trained weights are used, and we fine-tune the model on our dataset with a learning rate of 2e-5. For the Diffusion Policy, we employ DistilBERT to encode language instructions, adopting a methodology similar to YAY [44]. All models are trained for the same number of epochs, and only the final checkpoint is used for evaluation. To ensure a fair comparison, we avoid cherry-picking the checkpoint.
@@ 8.2 Efficient Inference
Fast inference speed is critical for deploying VLA models in real-world applications. However, as models grow in size and complexity, their inference speed tends to decrease significantly on server-grade hardware. To address this challenge and enhance model performance, we deploy our models on the vLLM framework [27]. This approach leverages optimized infrastructure to maximize inference efficiency. In Table 6, we present a comparative analysis of control frequencies for DiVLA-2B, DiVLA-7B, and OpenVLA (a 7B model). Our results demonstrate remarkable speed improvements: the DiVLA-2B model achieves an impressive 82 Hz on an A6000 GPU, showcasing exceptional performance. Similarly, DiVLA-7B achieves a control frequency of 42 Hz, which is 8 times faster than OpenVLA at the same model size. These findings underscore the effectiveness of our optimizations in maintaining scalability without sacrificing speed, paving the way for broader real-world applicability of VLA models.
@@
Notably, we avoid using action chunking techniques employed in other works, such as Data-Scaling-Laws [30] and $\pi_{0}$ [7], as we find that while action chunking can improve inference speed, it adversely impacts generalization performance. Another key observation is that our model experiences significant performance degradation when evaluated at lower bit precisions, such as 8-bit and 4-bit. This aligns with findings in OpenVLA, which highlight performance fluctuations introduced by quantization. These results suggest that current VLA models may require specifically designed quantization methods to maintain performance while achieving fast inference speeds at low precision.
@@ 8.4 DiVLA can do Visual-Question-Answering
Previous works, such as RT-2 [9] and ECoT [64], have suggested that co-training with vision-language data helps preserve basic conversational functionality for Vision-Language Action (VLA) data. In this section, we demonstrate that despite not being co-trained with vision-language data, our method retains its chatting capability and is proficient in performing common visual-question-answering (VQA) tasks. However, these works have neither provided concrete examples of chat interactions nor open-sourced their results, leaving it unclear to what extent conversational capabilities are retained in such models. We give a few examples in Table 9.

### v2

@@ 4 Experiments
Figure 5: Examples of visual variations, including randomly placed distractors, different backgrounds, and distracting lighting. DiVLA is robust to visual changes in different scenarios.
@@ 4.2 Real-World Multi-Task Learning
We begin with a standard setting in which the model is trained on multiple tasks and completes each task based on different user queries. We designed five tasks: object selection, flip the vertically placed pot, placing a cube into a designated box, placing a cup onto a plate, and placing a cube inside a box. Detailed descriptions of these tasks are provided in the Appendix. The experimental results can be found in Table 1. We compare our method to the Diffusion Policy (Chi et al., 2023), TinyVLA (Wen et al., 2024), Octo (Octo Model Team et al., 2024), and OpenVLA (Kim et al., ). Notice that both Octo and OpenVLA is pre-trained on OXE (O’Neill et al., 2023), which is 25 times larger than our pre-trained datasets.
@@ 4.3 End-to-End Sorting on Real Robot
We evaluate the capability of DiVLA in an industrial setting, where a robot is tasked with sorting items into designated sectors within a large box based on object category. Specifically, we categorize items into four classes: (1) toy cars, (2) knit gloves, (3) stuffed toys, and (4) hex keys. The language instruction provided is “Sort all items into corresponding areas”. A total of 500 trajectories are collected as training data. The task is considered successful only if the robot successfully grasps the object and places it in the correct sector. The experimental setup is illustrated in Figure 2.
@@
4.4 Behavior Analysis of Robot Foundation Model
Deep learning has demonstrated superior performance and generalization capabilities compared to traditional methods. Despite this, its nature as a “black box” algorithm often raises concerns regarding trustworthiness. The lack of transparency into the model’s decision-making process makes it difficult to understand its actions. This section explores applications of our reasoning injection module, designed to improve model interpretability by illuminating its internal reasoning and explaining why certain actions might fail.
How does the model generalize to new objects? We conduct experiments on sorting tasks. Specifically, we ask the model to identify and sort unseen objects, that are not in the train data. We place four previously unseen objects on the plate, a stuffed toy cat, a pair of green gloves, a dark toy car, and a screwdriver.
From this experiment, we observe several interesting findings. First, the model demonstrates the ability to analogize to new objects. For instance, while the model has not been trained on a screwdriver, it correctly categorizes it as a hex key due to their visual similarity. This indicates that the model can generalize to new objects, not by direct recognition, but by comparing the object’s semantic features with those of known objects. Second, the model can identify the color of some objects; for example, it categorizes the green glove as “green glove” and the stuffed toy cat as “brown toy cat.” Although these experiments are not exhaustive, they represent an important step towards the explainable generalization of robot models.
Failure case analysis via self-generated reasoning. A key advantage of our model is its ability to generate natural language rationales alongside its output actions, providing valuable insight into its decision-making process. By observing these reasoning phrases, we can understand what the model is “thinking” at each step. For instance, as illustrated in Figure 6, the model might initially identify a toy car and generate the reasoning, “grabbing the toy car,” indicating its intent to pick it up. However, if we intervene by placing a hex key in the gripper instead, the reasoning dynamically shifts to “grabbing the hex key.” This adaptive reasoning allows the model to adjust its subsequent actions, correctly sorting the hex key despite the unexpected change. This dynamic reasoning not only enhances the transparency and interpretability of the model’s actions but also contributes to its robustness. The integrated reasoning module allows for a form of self-correction, where the generated rationale guides and refines the action output, leading to more reliable performance even in the face of unexpected situations or perturbations. This capability suggests that the model is not simply executing pre-programmed actions, but rather engaging in a more flexible, context-aware decision-making process, akin to an internal dialogue.
Figure 6: What internal processes guide a model’s actions? We illustrate this using an example of DiVLA’s reasoning, inferred from shifts in its behavior based on changes to the target object. The model initially intends to grasp a “toy blue car”, but when presented with a hex key, its reasoning instantly redirects to sorting the “hex key” instead. This demonstrates one approach to interpreting a robot’s actions.
@@ 4.4 Behavior Analysis of Robot Foundation Model
Figure 8: Examples of various unseen objects in zero-shot bin-picking tasks. The unseen objects vary significantly in size.
@@ 4.6 Adapt to Real-World Bimanual Robot
Figure 9: (a) Environmental setup for the bimanual robot, (b) Table bussing setup, (c-f) All tableware and trash items used in the bussing table task evaluation.
In this section, we examine DiVLA’s adaptability to a dual-arm robot. Inspired by $\pi_{0}$ (Black et al., 2024), we designed a table bussing task that involves clearing table with various objects. This task was adapted for a bimanual robot setup: all tableware should be placed on a panel to the left, while trash items are dropped into a bin on the right. Similar to our factory sorting task, we evaluated the model’s performance using both seen objects and a combination of seen and unseen objects. The environment setup, along with all objects used for training and evaluation, is shown in Figure 9. Our evaluation consisted of twelve trials, with three to five objects randomly placed on the table. The success rate is computed by how many objects are correctly placed.
Our results show that our model successfully completes tasks in most cases when the object has appeared in the training data, achieving an average success rate of 72.9% on seen objects. In contrast, the Diffusion Policy and OpenVLA achieve 45.8% and 0% success rates. For tasks involving both seen and unseen objects, DiVLA achieves up to a 70.8% success rate, a slight decrease in success rate compared to the seen one, showing remarkable generalization to objects with different colors and shapes. Finally, DiVLA demonstrates the ability to recognize unseen objects, particularly by responding to object color. For example, it categorizes a Sprite can as a ”green can” and correctly places it in the trash bin. This observation further supports the idea that reasoning contributes to generalization.
@@ 5 Conclusion
In this work, we present DiVLA, a state-of-the-art vision-language-action model that delivers strong performance in both simulations and real-world scenarios, including single-arm and dual-arm robots. The core of our method lies in combining the next-token prediction objectives and diffusion models: the former for task reasoning and the latter for action prediction. We introduce a reasoning reuse module to enhance action generation. Through extensive evaluations in both simulations and across multiple real-world embodiments, we demonstrate that DiVLA outperforms several SOTA robot models. Additionally, we show that DiVLA has robust generalization capabilities, adapting effectively to new instructions, tasks, and environments. Our research offers a novel perspective on designing VLA models, encouraging a rethinking of how reasoning can be reused to facilitate end-to-end policy learning.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ 6.1 Evaluation Tasks
As described in Section 4, we evaluate our DiVLA and baselines on multi-task learning, factory sorting, zero-shot bin-picking, and table bussing. In this section, we provide details on these tasks. All of these tasks are depicted in Figure 10 and 11.
For multi-task learning, we evaluate all methods on 5 tasks for multi-task learning and 5 visual generalization tasks. We evaluate each method with a total of 77 trials for multi-task learning and 45 trials for visual generalization. We provide the number of demonstrations and the average length of these tasks in 3. A brief description of these tasks is as follows:
@@
Moreover, detailed descriptions for factory sorting, zero-shot bin-picking, and table bussing are provided in Sections 4.3, 4.5, and 4.6, respectively.
# Task # of Demonstrations Avg. Traj. Length
1 Object selection 160 91
2 Flip the vertically placed pot 120 137
3 Place cube inside box with closed lid 50 146
4 Place cup onto the plate 100 107
5 Place cube into yellow/blue Box 100 90
@@ 6.2 Detailed Results
Task $\setminus$ Method DP OpenVLA DiVLA-2B
Factory Sorting 0 0 60
@@ 7 Implementation Details
For the baselines, we follow a uniform training strategy to ensure consistency. For OpenVLA, the original implementation uses only a single camera view. Since our method use three camera views on both single arm Franka robot and bimanual AgileX robot, it would be unfair to compare the vanilla implementation of OpenVLA that uses only a single camera view. Therefore, for the purpose of fair comparison, we extend OpenVLA to utilize three camera views, processing each view through the same visual encoder and concatenating their outputs for further analysis. OpenVLA’s pre-trained weights are used, and we fine-tune the model on our dataset with a learning rate of 2e-5. Training is conducted over 20 epochs, as we find that OpenVLA typically requires longer training times for convergence. For the Diffusion Policy, we employ DistilBERT to encode language instructions, adopting a methodology similar to YAY (Shi et al., 2024). Table 7 shows the performance of the single-view OpenVLA baseline. In the sorting task, the success rate drops significantly from 45.3% with three views to just 12.7% with a single view, highlighting the importance of our proposed multi-view approach.
Table 7: Ablation study on OpenVLA using one camera view and three camera views. For a fair comparison, our main experiments evaluate OpenVLA using three camera views with existing methods. This section presents an ablation study comparing these results with those obtained using a single-camera view to demonstrate the benefits of our multi-view approach.
Method/Sorting Seen Mixed Cluttered Seen Cluttered Mixed Avg.
OpenVLA (3 views) 56.2 46.2 45.0 33.8 45.3
OpenVLA (1 view) 26.4 17.5 4.8 2.1 12.7
Table 8: Ablation study on reasoning injection module.
In-Distribution
Model $\setminus$ Tasks Task 1 Task 2 Task 3 Task 4 Task 5 Avg.
DiVLA-2B 100 100 63.6 63.6 90.9 83.6
w/o reasoning injection 66.7 66.7 45.5 45.5 27.3 50.3
@@ 8.2 Efficient Inference
Fast inference speed is critical for deploying VLA models in real-world applications. However, as models grow in size and complexity, their inference speed tends to decrease significantly on server-grade hardware. To address this challenge and enhance model performance, we deploy our models on the vLLM framework (Kwon et al., 2023). This approach leverages optimized infrastructure to maximize inference efficiency. In Table 5, we present a comparative analysis of control frequencies for DiVLA-2B, DiVLA-7B, and OpenVLA (a 7B model). Our results demonstrate remarkable speed improvements: the DiVLA-2B model achieves an impressive 82 Hz on an A6000 GPU, showcasing exceptional performance. Similarly, DiVLA-7B achieves a control frequency of 42 Hz, which is 8 times faster than OpenVLA at the same model size. These findings underscore the effectiveness of our optimizations in maintaining scalability without sacrificing speed, paving the way for broader real-world applicability of VLA models.
While vLLM accelerates our method, the improvement in inference time is less dramatic than reported for LLMs. Even without vLLM, DiVLA-2B and DiVLA-7B achieve respectable frame rates of 74Hz and 30Hz, respectively. Notably, this is still six times faster than OpenVLA, which has a similar model size (7 billion parameters).
@@
Another key observation is that our model experiences significant performance degradation when evaluated at lower bit precisions, such as 8-bit and 4-bit. This aligns with findings in OpenVLA, which highlight performance fluctuations introduced by quantization. These results suggest that current VLA models may require specifically designed quantization methods to maintain performance while achieving fast inference speeds at low precision.
Model $\setminus$ Tasks Following Novel Instruction on Four Tasks
Watermelon Watermelon $\rightarrow$ Blue Paper Trash $\rightarrow$ Lemonade
OpenVLA (Kim et al., ) 2/3 0/3
DiVLA-2B 2/3 2/3
Model $\setminus$ Tasks Red Pepper $\rightarrow$ Blue Cup Watermelon $\rightarrow$ Lemonade $\rightarrow$ Blue Paper Trash
OpenVLA (Kim et al., ) 0/3 0/3
DiVLA-2B 1/3 1/3
8.3 Following Novel Instruction
The previous section primarily evaluated the model’s performance on in-distribution task commands. In this section, we assess the model’s ability to follow novel instructions, specifically focusing on its generalization to previously unseen commands. We introduce new instructions that prompt the model to pick up unseen items and follow sequential instructions. Specifically, the instruction templates are “Pick up {obj}.” and “Pick up {obj} first, then pick up the {obj}, finally pick up {obj}.” We tested the model on four objects: 1) a watermelon, 2) a lemonade, 3) a blue paper trash, and 4) a red pepper. This is an extremely challenging task, as these novel instructions are absent from both the Droid dataset and our collected data. We evaluated four new instructions, with results summarized in Table 9.
Our findings indicate that both OpenVLA and DiVLA-2B can recognize these unseen objects and perform basic pick-and-place tasks. However, when it comes to complex sequential tasks, OpenVLA fails to interpret the instructions accurately, instead randomly selecting items. In contrast, our method correctly follows the instructions, picking up objects in the specified sequence. We hypothesize that by learning to decompose long tasks into subtasks, our method acquires a generalized ability to understand complex, multi-step instructions. While OpenVLA can execute simpler commands like “Pick up the watermelon,” it struggles with more advanced instructions requiring item selection in a specific order. Additionally, we observe a decrease in grasping precision when the model processes novel instructions, indicating that the novelty of instructions introduces further complexity to task execution.
8.4 Ablation Study on Reasoning Injection Module
Our core contribution is the proposed reasoning module, which enables the network to complete long-horizon tasks via direct prompting. This section presents an ablation study of this module, following the real-world experimental setup described in Section 4.2. We train a model using the same methodology, but without the reasoning injection module. Table 8 presents the results. We observe a significant performance drop compared to our baseline model when the reasoning injection module is removed. We hypothesize that the reasoning module facilitates task decomposition, resulting in more fine-grained action generation, which simplifies the learning process.
@@ 8.6 DiVLA can do Visual-Question-Answering
Previous works, such as RT-2 (Brohan et al., 2023) and ECoT (Zawalski et al., 2024), have suggested that co-training with vision-language data helps preserve basic conversational functionality for Vision-Language Action (VLA) data. In this section, we demonstrate that despite not being co-trained with vision-language data, our method retains its chatting capability and is proficient in performing common visual-question-answering (VQA) tasks. However, these works have neither provided concrete examples of chat interactions nor open-sourced their results, leaving it unclear to what extent conversational capabilities are retained in such models. We give a few examples in Table 11.